# MarS-FM: Generative Modeling of Molecular Dynamics via Markov State Models

### v1

@@ Abstract
11footnotetext: Work done as an intern at Valence Labs22footnotetext: Corresponding author: kacper.kapusniak@cs.ox.ac.uk
@@ 2 Preliminaries and Setting
A molecular system of $N$ atoms, in thermal equilibrium at temperature $T$, explores conformations according to the Boltzmann distribution $\mu$, i.e. $q\sim\mu(q)\propto\exp(-{\mathcal{H}}(q)/(k_{B}T))$ with $q=(x,\xi)\in{\mathcal{X}}$ denoting the system’s position and velocity, $\mathcal{H}$} the Hamiltonian, and $k_{B}$ the Boltzmann constant. We focus on proteins, whose motions—such as local unfolding or cryptic pocket formation—arise from structural fluctuations. To identify these conformational changes and quantify their frequencies (free energy), we aim to sample structures from the Boltzmann distribution $\mu$. This enables the measurement of observables $\phi$—functions defined on the phase space $\mathcal{X}$}—via expectations $\mathbb{E}_{x\sim\mu}[\phi(x)]$.
@@ 5.1 Preliminary investigation: Tetrapeptides
Torsions (bb) Torsions (sc) Torsions (all) TICA-0 TICA-0,1 joint MSM states Macrostate MAE MD (Oracle) 0.10 0.06 0.08 0.20 0.27 0.21 MDGen-1000 0.13 0.09 0.11 0.23 0.32 0.23 1.13 MDGen-200 0.14 0.10 0.12 0.24 0.33 0.27 1.12 MarS-FM $\oplus$ MDGen-200 (ours) 0.12 0.09 0.10 0.21 0.30 0.23 0.83 MarS-FM (ours) 0.11 0.09 0.10 0.21 0.29 0.22 0.63
@@ 5.2 MD-Cath: Proteins with sequence dissimilarity and large domain motions
Our main evaluation is based on the MD-Cath dataset from Mirarchi et al. (2024). This includes $5398$ domains up to $500$ residues (average number of residues $137$), simulated at $5$ different temperatures, spanning $5$ random replicas per setting. The average length of each trajectory is $464$ ns and frames are saved at $1$ ns intervals. The dataset is already built to ensure protein diversity by considering non-homologous domains at the $S20$ ($20\%$) homology level—further details are reported in Appendix B.
@@
Extended baselines. We benchmark different variants of MDGen to highlight how advantages of our framework cannot be simply replicated by changing the lag time or altering the sampling approach. As such, we train MDGen with window size $K=20$ and $K=100$, corresponding to a total lag time of $20$ ns and $100$ ns, respectively. Additionally, to demonstrate that the improvements of MarS-FM cannot be simply ascribed to a reduction of autoregressive calls, we also evaluate MDGen in parallel, meaning that we sample conformations only conditioned on the input frame. Finally, we also report the performance of one MD replica against 4 held-out ones (and average over all possible 5 combinations) to be treated as oracle performance.
@@ B.3.1 Tetrapeptides
Here, $G_{i}$ denotes the free energy of macrostate $i$, $k_{B}$ is the Boltzmann constant, $T$ is the temperature, and $\pi_{i}$ is the stationary distribution (normalized histogram count) of macrostate $i$. We use a small floor value (e.g., $10^{-4}$) to avoid numerical instability in the logarithm. Since we do not have access to additional replicas, we did not report the mMAE of the Oracle.
@@ B.4 Training Details
We follow the hyper-parameters of Jing et al. (2024b) as closely as possible. Throughout all experiments we use an exponential moving average with decay $0.999$, and AdamW (Loshchilov & Hutter, 2017) optimizer with a learning rate of $1\times 10^{-4}$. We use an architecture described in Appendix A.1 with 5 transformer layers, $384$-dimensional token embeddings, $16$-head multi-head attention, and an IPA layers with $4$ heads, $32$-dimensional head size, and $8$ query–key as well as $8$ value points.
For the tetrapeptide dataset we recreate the setup of Jing et al. (2024b), sampling MDGen-1000 from the released checkpoint and training our MarS-FM model for $1000$ epochs and MDGen-200 for $2500$ epochs, both with a batch size of $8$.
On MD-Cath dataset, MarS-FM is run for $1000$ epochs with batch size $8$ while sampling training pairs from two source clusters ($x_{0}$) and two corresponding destination clusters ($x_{1}$) with 12 sequences per cluster; MDGen-100 is trained for the same epoch budget with batch size $1$, and MDGen-20 for $2500$ epochs with batch size $4$.

### v2

@@ Abstract
11footnotetext: Work done as an intern at Valence Labs/Recursion22footnotetext: Corresponding author: kacper.kapusniak@cs.ox.ac.uk
@@ 2 Preliminaries and Setting
A molecular system of NN atoms, in thermal equilibrium at temperature $T$, explores conformations according to the Boltzmann distribution $\mu$, i.e. $q\sim\mu(q)\propto\exp(-{\mathcal{H}}(q)/(k_{B}T))$ with $q=(x,\xi)\in{\mathcal{X}}$ denoting the system’s position and velocity, $\mathcal{H}$} the Hamiltonian, and $k_{B}$ the Boltzmann constant. We focus on proteins, whose motions—such as local unfolding or cryptic pocket formation—arise from structural fluctuations. To identify these conformational changes and quantify their frequencies (free energy), we aim to sample structures from the Boltzmann distribution $\mu$. This enables the measurement of observables $\phi$—functions defined on the phase space $\mathcal{X}$}—via expectations $\mathbb{E}_{x\sim\mu}[\phi(x)]$.
@@ 5.1 Preliminary investigation: Tetrapeptides
Torsions ($b$) Torsions (sc) Torsions (all) TICA-0 TICA-0,1 joint MSM states Macrostate MAE MD (Oracle) 0.10 0.06 0.08 0.20 0.27 0.21 MDGen-1000 0.13 0.09 0.11 0.23 0.32 0.23 1.13 MDGen-200 0.14 0.10 0.12 0.24 0.33 0.27 1.12 MarS-FM $\oplus$ MDGen-200 (ours) 0.12 0.09 0.10 0.21 0.30 0.23 0.83 MarS-FM (ours) 0.11 0.09 0.10 0.21 0.29 0.22 0.63
@@ 5.2 MD-Cath: Proteins with sequence dissimilarity and large domain motions
Our main evaluation is based on the MD-Cath dataset from Mirarchi et al. (2024). This includes $5398$ domains up to $500$ residues (average number of residues $137$), simulated at 55 different temperatures, spanning 55 random replicas per setting. The average length of each trajectory is $464$ ns and frames are saved at 11 ns intervals. The dataset is already built to ensure protein diversity by considering non-homologous domains at the $S20$ ($20\%$) homology level—further details are reported in Appendix B.
@@
Extended baselines. We benchmark different variants of MDGen to highlight how advantages of our framework cannot be simply replicated by changing the lag time or altering the sampling approach. As such, we train MDGen with window size $K=20$ and $K=100$, corresponding to a total lag time of 2020 ns and $100$ ns, respectively. Additionally, to demonstrate that the improvements of MarS-FM cannot be simply ascribed to a reduction of autoregressive calls, we also evaluate MDGen in parallel, meaning that we sample conformations only conditioned on the input frame. Finally, we also report the performance of one MD replica against 4 held-out ones (and average over all possible 5 combinations) to be treated as oracle performance.
@@ B.3.1 Tetrapeptides
Here, $G_{i}$ denotes the free energy of macrostate $i$, $k_{B}$ is the Boltzmann constant, $T$ is the temperature, and $\pi_{i}$ is the stationary distribution (normalized histogram count) of macrostate ii. We use a small floor value (e.g., $10^{-4}$) to avoid numerical instability in the logarithm. Since we do not have access to additional replicas, we did not report the mMAE of the Oracle.
@@ B.4 Training Details
We follow the hyper-parameters of Jing et al. (2024b) as closely as possible. Throughout all experiments we use an exponential moving average with decay $0.999$, and AdamW (Loshchilov & Hutter, 2017) optimizer with a learning rate of $1\times 10^{-4}$. We use an architecture described in Appendix A.1 with 5 transformer layers, $384$-dimensional token embeddings, $16$-head multi-head attention, and an IPA layers with 44 heads, $32$-dimensional head size, and 88 query–key as well as 88 value points.
For the tetrapeptide dataset we recreate the setup of Jing et al. (2024b), sampling MDGen-1000 from the released checkpoint and training our MarS-FM model for $1000$ epochs and MDGen-200 for $2500$ epochs, both with a batch size of 88.
On MD-Cath dataset, MarS-FM is run for $1000$ epochs with batch size 88 while sampling training pairs from two source clusters ($x_{0}$) and two corresponding destination clusters ($x_{1}$) with 12 sequences per cluster; MDGen-100 is trained for the same epoch budget with batch size 11, and MDGen-20 for $2500$ epochs with batch size 44.